OpenStudy (anonymous):

In a certain county, the number of charter schools is 6 less than twice the number of alternative schools. We know that there are 52 charter schools in the county. How many alternative schools are in the county? (4 points) Select one: a. 58 b. 44 c. 29 d. 23 plz help me

OpenStudy (dumbcow):

C = 2A - 6 C = 52 --> 2A -6 = 52 solve for A

OpenStudy (dumbcow):

hmm (4 points) .... is this a test question, these online math tests are a joke and if you are gonna cheat why not google question or use yahoo answers

OpenStudy (anonymous):

is not a test :/ its just practice

OpenStudy (anonymous):

but thank you for your help i was lost and know i know the answer is 23

OpenStudy (dumbcow):

no way to know, its ok i see it a lot and really im frustrated with the system making it so easy to cheat check your algebra again , A is not 23

OpenStudy (anonymous):

D'': i m nut cheateing i m just need help

OpenStudy (dumbcow):

im helping, i set up equation now you need to isolate the variable 2A -6 = 52 ...

OpenStudy (anonymous):